# The pretty bad measurement

### v1

@@ Abstract
The quantum state discrimination problem has Alice sending a quantum state to Bob who wins if he correctly identifies the state. The pretty good measurement, also known as the square root measurement, performs pretty well at this task. We study the version of this problem where Bob tries to lose with the greatest probability possible (which is harder than it sounds). We define the pretty bad measurement which performs pretty well at this task, or in other words, pretty poorly for the original task. We show that both the pretty good measurement and the pretty bad measurement are always no worse than blind guessing at their respective tasks. As an application, we apply the pretty bad measurement to the quantum state anomaly detection problem and show how to avoid pretty bad qubits.
@@
Figure 1: Alice prepares $\rho_{i}\in\{\rho_{1},\ldots,\rho_{k}\}$ and sends it to Bob. Bob wins the game if he correctly identifies the index $i$ of the quantum state.
One can easily show that Bob can win this game perfectly if and only if the states have orthogonal supports, i.e., $\mathrm{Tr}(\rho_{i}\rho_{j})=0$ for all $i\neq j$. When this is not the case, Bob may wish to maximize his probability of winning. Suppose he measures with the POVM $(M_{1},\ldots,M_{k})$ and on outcome $j$ he guesses the state was $\rho_{j}$. Then, given state $\rho_{i}$, he guesses “I think the state is $\rho_{j}$” with probability $\mathrm{Tr}(\rho_{i}M_{j})$. In this case, he wins the game with probability
$\sum_{i=1}^{k}p_{i}\mathrm{Tr}(\rho_{i}M_{i}).$ (1)
If Bob wishes to maximize this probability then he is tasked with finding the best POVM $(M_{1},\dots,M_{k})$, also known as the minimum error measurement, and we define the best probability for Bob as
@@
Figure 2: Alice prepares $\rho_{i}\in\{\rho_{1},\ldots,\rho_{k}\}$ and sends it to Bob. Bob wins the game if he correctly excludes an index $j\neq i$ of the quantum state.
@@
If Bob wishes to find the measurement which best excludes a state, he wishes to minimize the probability of making an incorrect guess. More precisely, if Bob measures $\rho_{i}$ with the POVM $(M_{1},\ldots,M_{k})$, he makes an incorrect guess with probability $\mathrm{Tr}(\rho_{i}M_{i})$. Thus, he wishes to minimize
$\sum_{i=1}^{k}p_{i}\mathrm{Tr}(\rho_{i}M_{i}).$ (3)
@@
The pretty good measurement (PGM) was introduced in [16] as a way to approximate the best possible measurement. Sometimes PGM is optimal, some nice examples being for the hidden subgroup problem [4, 23], for the quantum change point problem [31], and in port-based teleportation [20]. Generally speaking, for problems with a large amount of symmetry, the PGM seems to perform well.
@@
To the best of our knowledge, comparing $\mathcal{P}_{\mathrm{PGM}}$ to the blind guessing probability, i.e., $1/k$, is not known except in trivial cases. This brings us to our first main result (which we prove in the appendix).
@@
For any set of states $\{\rho_{1},\ldots,\rho_{k}\}$ and any probability distribution $(p_{1},\ldots,p_{k})$, we have
$\mathcal{P}_{\mathrm{PGM}}\geq 1/k$ (8)
@@
Since $\alpha_{0}=\mathcal{P}_{\mathrm{PGM}}$, as this is the no-offset success probability, we have the following equation which follows directly from Eq. (11).
@@
By combining with Theorem 1, we have our main result, below, which is the counterpart to our PGM result, Theorem 1.
@@
For any set of states $\{\rho_{1},\ldots,\rho_{k}\}$ and any probability distribution $(p_{1},\ldots,p_{k})$, we have
$\mathcal{P}_{\mathrm{PBM}}\leq 1/k.$ (18)
@@
To summarize our discussion so far, we have
@@
recalling the definition of $P$ from Eq. (5). From this, we have
$\widetilde{P}=\sum_{i=1}^{k}\widetilde{p}_{i}\widetilde{\rho}_{i}=P.$ (22)
Denoting the PGM in the tilde setting as $(\widetilde{G}_{1},\ldots,\widetilde{G}_{k})$, we have
$\widetilde{G}_{i}=\widetilde{P}^{-1/2}\bigg{(}\widetilde{p}_{i}\widetilde{\rho}_{i}\bigg{)}\widetilde{P}^{-1/2}=B_{i}.$ (23)
Thus, one can view the PBM as the PGM if the problem instance is changed accordingly.
@@
Figure 3: If given a trine state, each with equal probability, we have $\mathcal{P}_{\mathrm{Best}}=\mathcal{P}_{\mathrm{PGM}}=2/3$, $\mathcal{P}_{\text{Random}}=1/3$ being the blind guessing probability, and $\mathcal{P}_{\mathrm{PBM}}=1/6$ approximating $\mathcal{P}_{\mathrm{Worst}}=0$.
@@
We have shown that the PGM always performs no worse than blind guessing and the PBM always performs no better than blind guessing. To get a better idea of how these two probabilities compare, we calculate them on many randomly generated examples. To generate random states, we can use the Bloch sphere to define a parameterized qubit $\rho(\theta,\phi)$ as
@@
where $X$, $Y$, and $Z$ are the Pauli operators and ${\theta,\phi\in(0,2\pi)}$ are real parameters which we randomly generate. We generate $1000$ random qubit instances $\{\rho_{1},\ldots,\rho_{k}\}$ for $k=2$ in FIG. 4 and for $k=6$ in FIG. 5, each with randomly generated probabilities $(p_{1},\ldots,p_{k})$ as well.
Figure 4: Computing PGM and PBM for $1000$ instances of discriminating between $2$ randomly chosen qubit states (with randomly chosen probability distributions as well). Notice that PGM always performs better than $1/2$ and PBM always performs worse than $1/2$ (although they sometimes get close).
Figure 5: Computing PGM and PBM for $1000$ instances of discriminating between $6$ randomly chosen qubit states (with randomly chosen probability distributions as well). We notice that $\mathcal{P}_{\mathrm{PGM}}$ seemingly clusters around $\frac{2}{k}$ and $\mathcal{P}_{\mathrm{PBM}}$ seemingly clusters around $\frac{k-2}{k(k-1)}$ for small values of $k$, with $k=6$ illustrated here.
@@
As an example, we consider the case where Alice promises to send to Bob $7$ copies of the state $\rho=|0\rangle\langle 0|$ but an anomaly occurs creating the state $\sigma=|\psi\rangle\langle\psi|$ where $|\psi\rangle=\gamma|0\rangle+\sqrt{1-\gamma^{2}}|1\rangle$ for some fixed parameter $\gamma\in[0,1]$ known to both Alice and Bob. FIG. 6 illustrates the probabilities $\mathcal{P}_{\mathrm{Best}}$, $\mathcal{P}_{\mathrm{PGM}}$, $\mathcal{P}_{\mathrm{PBM}}$, $\mathcal{P}_{\mathrm{Worst}}$ and the blind guessing probability $1/k$, as a function of $\gamma$, below.
Figure 6: Probabilities for the quantum state discrimination task defined by the states $|\psi_{i}\rangle=|0\rangle^{\otimes{i-1}}\otimes|\psi\rangle\otimes|0\rangle^{\otimes{7-i}}$, where $|\psi\rangle=\gamma|0\rangle+\sqrt{1-\gamma^{2}}|1\rangle$ and $i\in\{1,\ldots,7\}$. Observe that while it is hard to notice a difference between $\mathcal{P}_{\mathrm{PGM}}$ and the best probability, $\mathcal{P}_{\mathrm{PBM}}$ departs from the worst probability around $\gamma=0.4$. All of these probabilities coincide at $\gamma=1$ since all the states are the same in this case (see Eq. (28)).
@@ Conclusions.
In this work we introduced the pretty bad measurement as a means to approximate the worst guessing probability for quantum state discrimination and compared it to the pretty good measurement. We showed that they have an intricate relationship and prove that each compare favorably to simply randomly guessing. Finally, we examined the quantum anomaly detection/avoidance problem and gave an illustrative example showing how these measurements perform for detecting/avoiding parameterized quantum states.
@@ Acknowledgements.
This research was funded in part by the Commonwealth of Virginia’s Commonwealth Cyber Initiative (CCI) under grant number 467714.
@@ Appendix A Linear algebra background
Appendix B Proof of Theorem 1
@@
proving that $\mathcal{P}_{\mathrm{PGM}}\geq 1/k$, as desired.

### v4

@@ Abstract
The quantum state discrimination problem has Alice sending a quantum state to Bob who wins if he correctly identifies the state. The pretty good measurement, also known as the square root measurement, performs pretty well at this task. We study the version of this problem where Bob tries to lose with the greatest probability possible (which is harder than it sounds). We define the pretty bad measurement which performs pretty well at this task, or in other words, pretty poorly for the original task. We show that both the pretty good measurement and the pretty bad measurement are always no worse than blind guessing at their respective tasks. As an application, we apply the pretty bad measurement to the quantum state anomaly detection problem and show how to avoid pretty bad qubits. [^1] Published under the title “Quantum state exclusion through offset measurement” in Physical Review A.
@@
Figure 1: Alice prepares $\rho_{i}\in\{\rho_{1},\ldots,\rho_{k}\}$ and sends it to Bob. Bob wins the game if he correctly identifies the index $i$.
One can easily show that Bob can win this game perfectly if and only if the states have orthogonal supports, i.e., $\mathrm{Tr}(\rho_{i}\rho_{j})=0$ for all $i\neq j$. When this is not the case, Bob may wish to maximize his probability of winning. Suppose he measures with the POVM $(M_{1},\ldots,M_{k})$ and on outcome $j$ he guesses the state was $\rho_{j}$. Then, given state $\rho_{i}$, he guesses “I think the state is $\rho_{j}$” with probability $\mathrm{Tr}(\rho_{i}M_{j})$. In this case, he wins the game with probability $\sum_{i=1}^{k}p_{i}\mathrm{Tr}(\rho_{i}M_{i})$. If Bob wishes to maximize this probability, then he is tasked with finding the best POVM $(M_{1},\dots,M_{k})$, known as the minimum error measurement, and we define the best probability for Bob as
@@
Figure 2: Alice prepares $\rho_{i}\in\{\rho_{1},\ldots,\rho_{k}\}$ and sends it to Bob. Bob wins the game if he correctly excludes an index $j\neq i$.
@@
If Bob wishes to find the measurement which best excludes a state, he wishes to minimize the probability of making an incorrect guess. More precisely, if Bob measures $\rho_{i}$ with the POVM $(M_{1},\ldots,M_{k})$, he makes an incorrect guess with probability $\mathrm{Tr}(\rho_{i}M_{i})$. Thus, he wishes to minimize $\sum_{i=1}^{k}p_{i}\mathrm{Tr}(\rho_{i}M_{i})$.
@@
The pretty good measurement (PGM), first discussed in [5, 6, 19], is a way to approximate the best possible measurement. Sometimes PGM is optimal, some nice examples being for the hidden subgroup problem [4, 27], for the quantum change point problem [36], and in port-based teleportation [23]. Generally speaking, for problems with a large amount of symmetry, the PGM seems to perform well.
@@
We now present a theorem which compares $\mathcal{P}_{\mathrm{PGM}}$ to the blind guessing probability, i.e., $1/k$.
@@
[31] For any set of states $\{\rho_{1},\ldots,\rho_{k}\}$ and probability distribution $(p_{1},\ldots,p_{k})$, we have
$\mathcal{P}_{\mathrm{PGM}}\geq\frac{1}{k}+\frac{(1-k\mathcal{P}_{\mathrm{Best}})^{2}}{k(k-1)}$ (6)
@@
We prove a slightly weaker version of this theorem in the appendix using the concept of offsets, discussed shortly.
@@
Since $\alpha_{0}=\mathcal{P}_{\mathrm{PGM}}$, as this is the no-offset success probability, we have the following equation which follows directly from Eq. (9)
@@
By combining with Theorem 1, we have our main result, below, which is the counterpart to the PGM bound, Theorem 1.
@@
For any set of states $\{\rho_{1},\ldots,\rho_{k}\}$ and probability distribution $(p_{1},\ldots,p_{k})$, we have
$\mathcal{P}_{\mathrm{PBM}}\leq\frac{1}{k}-\frac{(1-k\mathcal{P}_{\mathrm{Best}})^{2}}{k(k-1)^{2}}.$ (16)
@@
As a consequence of our discussion so far, we have
@@
recalling the definition of $P$ from Eq. (3). We also have that $\widetilde{P}=\sum_{i=1}^{k}\widetilde{p}_{i}\widetilde{\rho}_{i}=P$. Denoting the PGM in the tilde setting as $(\widetilde{G}_{1},\ldots,\widetilde{G}_{k})$, we can calculate that $\widetilde{G}_{i}=\widetilde{P}^{-1/2}\bigg{(}\widetilde{p}_{i}\widetilde{\rho}_{i}\bigg{)}\widetilde{P}^{-1/2}=B_{i}$. Thus, one can view the PBM as the PGM if the problem instance is changed accordingly.
Remark: We note that the mixing of POVM elements considered above is studied more generally in the context of reverse channels and reverse observables which have been used in the study of compatibility [15].
@@
Figure 3: If given a trine state, each with equal probability, we have $\mathcal{P}_{\mathrm{Best}}=\mathcal{P}_{\mathrm{PGM}}=2/3$, $\mathcal{P}_{\text{Random}}=1/3$ being the blind guessing probability, and $\mathcal{P}_{\mathrm{PBM}}=1/6$ approximating $\mathcal{P}_{\mathrm{Worst}}=0$. The pretty good measurement is given by the POVM $G_{i}=\frac{2}{3}|\psi_{i}\rangle\langle\psi_{i}|$, the pretty bad measurement is given by the POVM $B_{i}=\frac{1}{3}\sum_{j\neq i}|\psi_{j}\rangle\langle\psi_{j}|$, and the worst measurement is given by $W_{i}=\frac{2}{3}|\psi_{i}^{\perp}\rangle\langle\psi_{i}^{\perp}|$, where $|\psi_{i}^{\perp}\rangle\langle\psi_{i}^{\perp}|$ is a $\pi/2$-rotation of $|\psi_{i}\rangle$.
@@
We have discussed how the PGM always performs no worse than blind guessing and consequently proved that the PBM always performs no better than blind guessing. To get a better idea of how these two probabilities compare, we calculate them on many randomly generated examples. To generate random states, we can use the Bloch sphere to define a parameterized qubit $\rho(\theta,\phi)$ as
@@
where $X$, $Y$, and $Z$ are the Pauli operators and ${\theta,\phi\in(0,2\pi)}$ are real parameters which we randomly generate. We generate $1000$ random qubit instances $\{\rho_{1},\ldots,\rho_{k}\}$ for $k=2$ in FIG. 4 and for $k=6$ in FIG. 5, each with randomly generated probabilities $(p_{1},\ldots,p_{k})$.
Figure 4: Computing PGM and PBM for $1000$ instances of $2$ randomly chosen qubits (also with randomly chosen probabilities). Notice that PGM always performs better than $1/2$ and PBM always performs worse than $1/2$.
Figure 5: Computing PGM and PBM for $1000$ instances of $6$ randomly chosen qubits (also with randomly chosen probabilities). We notice that $\mathcal{P}_{\mathrm{PGM}}$ seemingly clusters around $\frac{2}{k}$ and $\mathcal{P}_{\mathrm{PBM}}$ seemingly clusters around $\frac{k-2}{k(k-1)}$ for small values of $k$, with $k=6$ illustrated here.
@@
As an example, we consider the case where Alice promises to send to Bob $7$ copies of the state $\rho=|0\rangle\langle 0|$ but an anomaly occurs creating the state $\sigma=|\psi\rangle\langle\psi|$ where $|\psi\rangle=\gamma|0\rangle+\sqrt{1-\gamma^{2}}|1\rangle$ for some fixed parameter $\gamma\in[0,1]$ known to both Alice and Bob. FIG. 6 illustrates the probabilities $\mathcal{P}_{\mathrm{Best}}$, $\mathcal{P}_{\mathrm{PGM}}$, $\mathcal{P}_{\mathrm{PBM}}$, $\mathcal{P}_{\mathrm{Worst}}$ and the blind guessing probability $1/k$, as a function of $\gamma$, below. For anomaly detection, [24, 38] showed that PGM is indeed the optimal measurement. Moreover they provide a closed-form expression for the optimal success probability. Therefore, using Eq. (15) we also have an expression for $\mathcal{P}_{\mathrm{PBM}}$.
Figure 6: Probabilities for the quantum state discrimination task defined by the states $|\psi_{i}\rangle=|0\rangle^{\otimes{i-1}}\otimes|\psi\rangle\otimes|0\rangle^{\otimes{7-i}}$, where $|\psi\rangle=\gamma|0\rangle+\sqrt{1-\gamma^{2}}|1\rangle$ and $i\in\{1,\ldots,7\}$. Observe that while $\mathcal{P}_{\mathrm{PGM}}=\mathcal{P}_{\mathrm{Best}}$ for all values of $\gamma$, $\mathcal{P}_{\mathrm{PBM}}$ departs from the worst probability around $\gamma=0.4$. All of these probabilities coincide at $\gamma=1$ since all the states are the same in this case (see Eq. (24)).
@@ Conclusions.
In this work we introduced the pretty bad measurement as a means to approximate the worst guessing probability for quantum state discrimination and compared it to the pretty good measurement. We showed that they have an intricate relationship and proved that each compare favorably to simply randomly guessing. Finally, we examined the quantum anomaly detection/avoidance problem and gave an illustrative example showing how these measurements perform for detecting/avoiding parameterized quantum states.
@@ Acknowledgements.
The authors thank Iman Marvian, Vincent Russo, Nathaniel Johnston, and Santiago Llorens for useful discussions. This research was funded in part by the Commonwealth of Virginia’s Commonwealth Cyber Initiative (CCI) under grant number 467714.
@@ Appendix A Linear algebra background
Appendix B Proof of a slightly weaker version of Theorem 1
@@
proving that $\mathcal{P}_{\mathrm{PGM}}\geq 1/k$.